4. You have a lung model whose airway flow is 90 liters/minute. The pressure at one end of the airway is 40 cm H2O, and the pressure at the other end of the airway is 5 cm H2O. What is the value for airway resistance in this model (watch your units!)?
Answer #1

Airway resistance is the ratio of change in pressure to flow rate of air.

=Change in airway pressure/ flow f air

Airway flow =90 L/minute= 1.5 L/sec

Pressure at one end of airway= 40 Cm of H2O

Pressure at other end of airway= 5 Cm of H2O

Change in airway pressure= 40-5= 35 Cm of H2O

Airway resistance = Change in airway pressure / Flow of air

   = 35/1.5= 23.3 Cm of H2O / sec